A rectangle's length is 11 inches longer than its width. The perimeter is 262 inches. Find the length and width.
Answer by algebriac(10)

Let the width of rectangle be w inches
Since rectangle's length is 11 inches longer than its width
Hence Length of rectangle= w+11 inches
Perimeter of rectangle =262 inches
Perimeter= 2(l+w)
2(w+11+w)=262
2w+11=131
2w=120
W=60

Hence Width= 60 inches
Length= 60+11=71 inches